Circular No. 8959

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 N1 (HOLMES) An apparently asteroidal object discovered by Robert E. Holmes, Jr.,

Charleston, IL, U.S.A., on CCD images taken with a 0.40-m f/5.8 Schmidt-Cassegrain reflector (discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance by other CCD astrometrists. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector) reports that the object appears slightly (∼ 10 percent) larger and more diffuse than nearby stars of similar brightness in a 23-min exposure from July 4.08 UT, while a 400-s exposure shows the object to be “soft” with very little central condensation and a diameter of 4′′, with a possible faint elongation 5′′ long in p.a. 225^{o} (but not significantly different from stellar images). L. Buzzi (Varese, Italy, 0.60-m f/4.6 reflector) writes that two series of stacked image totalling 30 min of exposure from July 5.03–5.07 show a nearly stellar coma apparently elongated ∼ 10′′ toward p.a. ∼ 190^{o}; his stacked exposures from July 6.0 in very good seeing show an image slightly larger than nearby stars of similar brightness but with no tail or elongation. S. Foglia reports that analysis of Holmes’ images from July 2.35, 5.4, and 6.4 show a stellar appearance, though the size of the images appear slightly larger than stars of similar brightness. J. Lacruz, La Cañada, Spain, reports that his “blinking” of sets of images taken (apparently with a 0.3-m reflector) on July 5.06 show a diffuse coma ∼ 12′′ in diameter, elongated to the southwest.

2008 UT  α_{2000}  δ_{2000}  Mag.
July 1.32643  23^{h}44^{m}50.^{s}01  +14^{o}16′12′′.3  20.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-N19.

T = 2009 Sept.24.954 TT  ω = 100.^{o}882
Ω = 357.474 }2000.0
q = 2.77594 AU  i = 115.556

COMET C/2008 J1 (BOATTINI) Visual total-magnitude estimates: May 12.11 UT, 12.3 (J. J. Gonzalez,

Leon, Spain, 20-cm reflector); June 1.92, 11.3 (M. L. Paradowski, Dabrowa, Poland, 25-cm refl.); 14.97, 10.8 (Paradowski); 27.98, 10.3 (Gonzalez); July 2.01, 10.8 (M. Meyer, Limburg, Germany, 25-cm refl.); 4.99, 10.7 (Meyer).

2008 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green